OpenStudy (anonymous):

Anthony is making a collage for his art class by picking shapes randomly. He has five squares, two triangles, two ovals, and four circles. Find P(a triangle or a square is chosen first).

OpenStudy (kropot72):

The total number of shapes of 4 kinds is 5 + 2 + 2 + 4 = 13 The total number of triangles and squares is 5 + 2 = 7 \[P(\triangle\ or\ \square\ chosen\ first)=\frac{total\ number\ of\ triangles\ +\ squares}{total\ number\ of\ all\ shapes}\]